OpenStudy (anonymous):

y''+y'-2=2t find the general solution of the nonhomogenous differential equation

OpenStudy (amistre64):

-2? or -2y?

OpenStudy (amistre64):

the homogenous part will be drastic enough; but the particuylar is pretty basic you have a linear so decipher it with: yp = At+b

OpenStudy (anonymous):

2t

OpenStudy (anonymous):

oh its -2y

OpenStudy (anonymous):

my bad

OpenStudy (amistre64):

im going to forgo the capital stuff, its hard enough for me to keep track of the lowercases :) yp = at+b yp' = a yp'' =0 yp''+ yp'-2yp = 2t

OpenStudy (amistre64):

a -2at-2b = 2t -2a = 2 ; a = -1 a-2b = 0 -1 -2b - 0 ; b=-1

OpenStudy (amistre64):

y = yh -t-1

OpenStudy (anonymous):

thanks